# Natural Language Generation from Visual Sequences: Challenges and Future Directions

### v1

@@ Abstract
The ability to use natural language to talk about visual content is at the core of human intelligence and a crucial feature of any artificial intelligence system. Various studies have focused on generating text for single images. In contrast, comparatively little attention has been paid to exhaustively analyzing and advancing work on multiple-image vision-to-text settings. In this position paper, we claim that any task dealing with temporally ordered sequences of multiple images or frames is an instance of a broader, more general problem involving the understanding of intricate relationships between the visual content and the corresponding text. We comprehensively analyze five tasks that are instances of this problem and argue that they pose a common set of challenges and share similarities in terms of modeling and evaluation approaches. Based on the insights from these various aspects and stages of multi-image-to-text generation, we highlight several open questions and suggest future research directions. We believe that these directions can advance the understanding of complex phenomena in this domain and the development of better models.
@@ 1 Introduction
Over the years, research in natural language generation has demonstrated the importance of grounding language in the visual modality to improve understanding and reasoning capabilities of models Baroni (2016); Beinborn et al. (2018); Wang et al. (2024b). Earlier work on visually conditioned language generation primarily focused on single-image-to-text tasks such as image captioning and visual question answering. However, many practical real-world vision-to-language (V2L) applications in several domains such as surveillance and media content creation require understanding and reasoning across multiple temporally ordered images or video frames. With the increase in availability of video and image sequence data, various tasks have been proposed over the past few years to develop and evaluate models that can generate text grounded in multiple images or frames of videos. While some proposed multi-image-to-text tasks such as Video Captioning (VC) Yao et al. (2015) and Multi-image/Video Question Answering (VIDQA) Zeng et al. (2017); Bansal et al. (2020) are reminiscent of popular single-image-to-text settings, other tasks such as Change Captioning (CC) Jhamtani and Berg-Kirkpatrick (2018) are unique both in terms of their objectives and the type of input-output data.
Task Input Output
@@
(a) Between consecutive images or video frames of input.
(b) Between consecutive sentences in ground-truth text.
Figure 1: Similarity scores obtained for the tasks along the visual and textual dimensions. We exclude Video Question Answering task from textual consistency analysis due to the lack of multi-sentence datasets.
Nevertheless, all multi-image-to-text tasks require models to reason along the temporal dimension of the visual input for generating the textual output. Therefore, in this position paper, we consider all multi-image-to-text tasks as instances of the broader problem of generating natural language output given a sequence of multiple temporally ordered images or video frames as input. Table 1 outlines a set of tasks that we consider as good representatives of this problem and presents the differences between the corresponding input and output data type. For all the tasks we provide examples in Table 2.
In summary, our main contributions are:
  • i.
    We quantitatively characterize each of the multi-image V2L tasks along two dimensions based on the complexities of their input-output data. We argue that the degree to which they vary along these dimensions is dependent on the corresponding task definitions and objectives.
  • ii.
    We highlight and discuss a common set of challenges relevant for multi-image-to-text generation such as accurate tracking and grounding of entities, ensuring coherence between output text segments, inter alia.
  • iii.
    We comprehensively review the evolution of modeling approaches, learning procedures, and evaluation protocols and provide a unified overview, which we believe will be useful for facilitating further advancements.
  • iv.
    Finally, we propose future research directions to improve the systems and methods used for tackling these tasks at different stages of the process.
2 Dimensions of Variation
While our focus is on the multi-image-to-text-problem, different tasks within this problem space may have different characteristics. Two relevant dimensions are the complexity of the visual input and of the textual output. For example, some of these tasks require models to generate succinct answers and descriptions, and others require generation of long-form textual narratives intended to also complement the visual information. These dimensions of variation across tasks typically tend to be crucial factors in making design choices with regard to developing model architectures and learning procedures. In terms of the visual input, depending on the objective of the task, images within the input sequence of each data sample could be comparable to each other or vary drastically to the point of being completely heterogeneous. For instance, in the CC task, where the goal is to localize and describe changes between a pair of images obtained from real-time surveillance cameras or large-scale remote-sensing snapshots, we hypothesize that the similarity between input images would be generally high. On the other hand, in tasks such as Visual Storytelling (VST) Huang et al. (2016), in which the input sequences typically depict an overarching narrative, we hypothesize low similarity between consecutive images within each data sample (<visual sequence, text> pair). Regarding textual output, we similarly posit that the consistency of consecutive sentences within each data sample could be high or low depending on the corresponding task objective.
To preliminarily test our intuitions, we quantitatively analyze a few datasets corresponding to each of these tasks. Specifically, for each data sample, we compute visual similarity scores between CLIP Radford et al. (2021) visual encoder embeddings of consecutive images in the input sequence and report the average score. In the same manner, we compute textual consistency scores between CLIP text encoder embeddings of consecutive sentences in the corresponding ground-truth text of each data sample. For this study, we randomly select 100 instances per task from five datasets—Spot-the-diff Jhamtani and Berg-Kirkpatrick (2018) for CC, VIST Huang et al. (2016) for VST, Charades Sigurdsson et al. (2016) for VC, MSVD-QA Xu et al. (2017) for VIDQA, and MAD-v1 Soldan et al. (2022) for Movie Auto Audio Description (Auto AD) Han et al. (2023b) (MAAD). [^1] Further details are provided in Appendix A. Figure 1 shows the distributions of similarity scores obtained for each of the tasks along the visual and textual dimensions. In terms of visual similarity, we observe that CC and VST indeed obtain maximum and minimum scores respectively, with other tasks ranging in between, confirming our intuitions. In terms of textual consistency, the differentiation is less evident. We observe that for the MAAD and VC tasks, consecutive sentences in the ground-truth text are relatively less consistent with each other. Using the average similarity scores across all data samples, we categorize the five tasks by placing them at different positions in the shared space between textual consistency and visual similarity (see Figure 2). Besides the two axes considered for this analysis, we note that tasks could also be compared along various other dimensions. We posit that this kind of analysis would serve as a meaningful guide for making modeling and evaluation decisions both for current and for novel future tasks in the multi-image-to-text landscape.
Figure 2: Tasks positioned in the visual-textual shared space using the similarity scores obtained for corresponding datasets. Interpretation: a task in the top right corner of the plot is both inconsistent at the textual level and dissimilar at the visual level.
3 Common Challenges
Multi-image-to-text tasks present a unique set of challenges. Some of these challenges are task-specific, but most of them are common across the five tasks. In this section, we describe these challenges and discuss how they instantiate for each of the tasks.
Entity tracking.
Identifying and tracking entities is an important requisite for accurately interpreting actions and relationships between them. This has been underlined by several works in both language understanding Paperno et al. (2016); Kim and Schuster (2023) and computer vision domains Narayan et al. (2017); Luo et al. (2021). In multi-image V2L tasks, the availability of input signals from two modalities makes the aspect of disambiguating entities more challenging. For instance, in the MAAD or VST tasks where the visual input is heterogeneous, entities in the input image sequences/video clips tend to ‘disappear’ in some of the images/frames at the intermediate temporal positions, while still being actively referenced in the textual input at the corresponding positions. Surikuchi et al. (2023) have denoted such cases as being temporally misaligned. To track entities accurately under such temporal misalignment, it is necessary to not only learn causal relations between people and objects in each of the modalities at corresponding positions, but also to obtain a cross-modal cross-temporal representation of all the relationships relevant to the scene/narrative. For the CC, VC, or VIDQA tasks, in which the input images/video frames are typically similar to each other, it is crucial to differentiate meaningful semantic entities and their changes from various distractions. While in CC, viewpoint changes or illuminations are considered as distractors and discarded, in the VIDQA task, entities relevant to answering the question need to be differentiated from others for accurate tracking. Effective tracking of entities would therefore require accounting for changes in appearance (including disappearance), capturing interactions, and correctly identifying occlusions.
Visual grounding.
Humans acquire language understanding through perception and interaction with the environment Barsalou (2008); Iverson and Goldin-Meadow (2005) and consequently this enables them to seamlessly ground language in visual data. Over the years, a great deal of work has been proposed to adapt the architecture and learning process of vision-language models for acquiring visual grounding Suglia et al. (2024). However, there are still significant challenges for achieving human-levels of grounding using computational models and this becomes more apparent when looked at from the perspective of multi-image-to-text tasks. For instance, in the VST or MAAD tasks, since language is typically ‘inconsistent’ (see Figure 2), it is also inadvertently under-specified semantically Pezzelle (2023) (e.g., ‘The boy on the bridge was waving to the tourists near the waterfall. A photographer over there clicks…’). Stories also tend to contain many abstract adverbs such as ‘often’ or ‘today’ and it has been shown that vision-language models struggle to disambiguate such under-specified text and accurately map phrases to regions in the image sequences/videos. Moreover, the amount of language informativeness—degree of information required for identifying the correct object Coppock et al. (2020)—could be inadequate in tasks such as VIDQA, particularly with the presence of confounding entities in various frames (e.g., input video of a football match and a question: ‘What is the color of the card the referee is holding?’). Also, when grounding objects, it has been shown that models often struggle to reliably capture spatial relationships Kamath et al. (2023). To summarize, beyond being merely descriptive, language in some multi-image-to-text tasks could also be complementary to the data in images/videos, making visual grounding challenging without access to relevant additional external knowledge.
Knowledge integration.
For some multi-image-to-text tasks, models would be required to utilize additional information beyond what is available in the input data. In VC or VIDQA tasks pertaining to certain domains such as news, the input video might not contain all the aspects needed to correctly describe its contents or answer questions about it Whitehead et al. (2018); Jin et al. (2023). To address this, various approaches often rely on using large pre-trained general purpose models or external knowledge bases such as ConceptNet Speer et al. (2017) to retrieve both factual and commonsense information. This method is commonly referred as retrieval-augmented generation (RAG) Lewis et al. (2020). Besides being sources for missing information, external knowledge bases are also leveraged for enriching the generated text with social or cultural contexts. For instance, in VST, some approaches use recognized visual objects in the input to retrieve concepts from external knowledge graphs for generating more engaging and figurative stories (e.g., concept of ‘graduation ceremony’ following the detection of an ‘academic gown’ object).
From knowledge selection/retrieval stage to accurately representing and utilizing it during text generation, the process of integrating external knowledge has various challenges. Robust retrieval systems are required which can holistically extract the essence of image sequences/video frames, including the various entities and their interrelationships. Typically, the retrieved knowledge is concatenated with input representations which are then used for generating text either through fine-tuning Yang et al. (2024) or by prompting general-purpose VLMs Bhattacharyya et al. (2023). However, this approach might lead to models either over or under utilizing the retrieved knowledge potentially leading to incoherent text Gao et al. (2023). To address this, we argue that fusion mechanisms which can effectively balance information from representations of both input data and the retrieved knowledge need to be developed. Furthermore, retrieving relevant knowledge from increasingly large databases could be computationally expensive, especially in the multi-image scenario. Ways to optimize retrieval components for improving efficiency is an active research area.
Textual coherence.
Coherence is the property of text that refers to the ordering of its constituents (words/sentences) and the way in which they relate to each other Althaus et al. (2004). Coherent text should have a consistent logical structure in which the events, interactions, and relationships between various elements are ordered in a meaningful way. It is an important aspect of discourse and has been studied extensively in neural language generation Pishdad et al. (2020). For several multi-image-to-text tasks, particularly the ones in which the output tends to be less ‘consistent’ along the text axis in Figure 2, it is challenging to ensure that multiple sentences in the generated output are locally coherent. In MAAD and VST tasks, where there are multiple characters and various interactions developing across the sequence of images/video frames, it is often difficult for models to balance between selecting the visual information and representing it in a cohesive/connected language Lei et al. (2020). This challenge is more apparent in the VST task, in which models are expected to keep track of multiple things such as emotional arcs or motivations of the characters and the overarching narrative Surikuchi et al. (2024). There is increasing work in unimodal text-only storytelling suggesting how using concepts of narratology Piper et al. (2021); Antoniak et al. (2024) such as Genette (1980)’s triangle can potentially aid models in generating stories with engaging and coherent structures. However, it still unclear how these theories can be applied to multimodal scenarios where the generated text needs to be consistent with image sequences/videos.
Theory of mind.
Theory of Mind (ToM), which is considered the basis of human social cognition Premack and Woodruff (1978), is described as the ability to understand and make inferences about the mental states (e.g., beliefs, intentions, and desires) of other people or living beings. In the context of multi-image V2L tasks, ToM refers to the ability of models to go beyond merely recognizing objects/actions and to reason about the mental states of entities depicted in the image sequences/videos. Although most of the tasks we consider in this work, besides VIDQA Mao et al. (2024), do not explicitly require ToM abilities, the skill is still relevant for all the tasks and closely connected to the other challenges and abilities discussed so far. For instance, in tasks such as VST, to causally connect heterogeneous images in the input sequence, models need to be equipped with different reasoning abilities pertaining to emotions and social perceptions/intentions. This enables generation of stories that reflect actions and mental states of the characters beyond literal interpretation of the visual data.
Recently, several ToM benchmarks have been proposed to assess general-purpose VLMs Gao et al. (2024); Chen et al. (2024b) along different aspects such as temporal localization of emotions, intentionality-understanding, and perspective-taking. However, these studies find that only models that are fine-tuned on curated ToM datasets exhibit any reasoning abilities, albeit not aligning with the well-established ToM theories/frameworks explaining human social cognition. Such curated data is scarcely available and it is unclear what alternative architectures or training objectives would enable models to obtain the ToM abilities required for multi-image V2L tasks.
4 Models Architectures
Figure 3: Outline of the architecture common across modeling approaches for multi-image-to-text tasks.
@@ 4 Models Architectures
Modeling approaches to multi-image-to-text tasks have evolved over time from being recurrent neural network (RNN)-based Hochreiter and Schmidhuber (1997) to being transformer-based Vaswani et al. (2017). More recent models directly leverage pre-trained large (vision)-language models (LLMs/VLMs), often in a zero-shot manner. In this section, we discuss this evolution and summarize the various state-of-the-art model architectures proposed for the five multi-image-to-text tasks. Architectures proposed for these tasks primarily comprise three modules—a vision encoder, a language decoder, and an intermediate module (typically referred as the projector/adapter) for adapting visual information into contextualized representations for text generation. We describe the functionality of these modules and review the design principles common across all the tasks in the proposed approaches. Furthermore, we also discuss how off-the-shelf pre-trained VLMs are currently being used to handle various multi-image-to-text tasks. Table 3 outlines a summary of the models and details related to the selection procedure are provided in Appendix B.
@@ 4.1 Vision Encoder
The primary purpose of a vision encoder in vision-to-language tasks is to extract information from the input visual sequence and to optimally encode it into a contextual representation that guides language generation. To achieve this, encoders in the proposed models follow multiple steps, some of which are common across the five multi-image-to-text tasks. First, a pre-trained vision model is utilized for extracting feature representations of the raw input sequences of images/video frames. Earlier approaches used convolutional neural network (CNN)-based vision models such as ResNet He et al. (2016) or R3D Tran et al. (2018) that are primarily pre-trained on the object detection task using large amounts of image/video data. Most of the recent models across the tasks use transformer-based vision models pre-trained for various image-only and image-text alignment objectives, e.g., CLIP-ViT-L Radford et al. (2021).
We note that besides the primary input sequence of images/video frames, models proposed for some of the tasks, e.g., MAAD, utilize additional input data such as close-ups of characters in the movie clips (exemplars) Han et al. (2023a). Furthermore, the TAPM Yu et al. (2021) model proposed for the VST task utilizes FasterRCNN Ren et al. (2015) for extracting such local character/object-level features alongside the global image-level features from ResNet. Following the extraction of visual features using pre-trained vision models, most vision encoders comprise an internal sequence-encoder component for learning relationships and dependencies between the individual image/frame-level features at different temporal positions. Some models implement this step either using RNNs or a transformer network with multi-head self-attention for learning temporal relationships and position-encoding mechanism for tracking the order of entities/events in the sequence.
Beyond these steps that are common across the tasks, vision-encoders may also contain additional task-specific steps for capturing the visual information in a way that suits the task’s objective better. For instance, the ViLA model Wang et al. (2025) for the VIDQA task utilizes a learnable Frame-Sampler component to efficiently select a small subset of frames that are most likely to contain the relevant information needed to answer the question. Another example with a task-specific step is the MSCM+BART model Hsu et al. (2020) for VST, in which the initial set of image objects/‘concepts’ are expanded using an external knowledge graph for generating diverse and informative stories. Despite these task-specific steps, we found that the vision encoder module in the architectures proposed for the various multi-image-to-text tasks share a common set of components that are broadly outlined in Figure 3.

### v2

# Natural Language Generation from Visual Events: Challenges and Future Directions

@@ Abstract
The ability to use natural language to talk about visual events is at the core of human intelligence and a crucial feature of any artificial intelligence system. In recent years, a substantial body of work in visually grounded NLP has focused on describing content depicted in single images. By contrast, comparatively less attention has been devoted to exhaustively modeling scenarios in which natural language is employed to interpret and talk about events presented through videos or sequences of images. In this position paper, we argue that any NLG task dealing with sequences of images or frames is an instance of the broader, more general problem of modeling the intricate relationships between visual events unfolding over time and the features of the language used to interpret, describe, or narrate them. Therefore, solving these tasks requires models to be capable of identifying and managing such intricacies. We consider five seemingly different tasks, which we argue are compelling instances of this broader multimodal problem. Consistently, we claim that these tasks pose a common set of challenges and share similarities in terms of modeling and evaluation approaches. Building on this perspective, we identify key open questions and propose several research directions for future investigation. We claim that improving language-and-vision models’ understanding of visual events is both timely and essential, given their growing applications. Additionally, this challenge offers significant scientific insight, advancing model development through principles of human cognition and language use.
@@ 1 Introduction
The ability to reason over visual events and describe them in natural language is a cornerstone of human communication and a crucial capability for intelligent systems operating in dynamic multimodal environments. From assistive technologies that interpret the world for visually impaired users to AI agents that must understand and narrate video content or interact with humans in real time, the demand for systems that can bridge visual perception and language over a temporal span is growing rapidly.
To date, much of the progress in visually grounded NLP has focused on language generation from a static image. Tasks such as image captioning Vinyals et al. (2015) and visual question answering Antol et al. (2015) have demonstrated the potential of multimodal models to ground linguistic expressions in visual input Baroni (2016); Beinborn, Botschen, and Gurevych (2018); Wang et al. (2024). However, the world is not static. Real-world applications such as media content creation, real-time sports commentary and news reporting, and interactive multimodal assistants frequently require reasoning about events as they unfold over time. This requires capturing key aspects of event understanding, including temporal dynamics, causality, and narrative structure, that arise from complex multimodal interaction.
In this position paper, we argue for a unified perspective: that all NLP tasks dealing with visual events, whether presented through videos or ordered image sets, are instances of a broader, more general problem. At the heart of this problem lies the need to model the intricate relationship between temporally ordered visual events and the structure, content, and function of the language used to interpret, describe, or narrate them. We claim that solving these tasks effectively requires models to be capable of identifying and managing such intricacies, leveraging the temporal, causal, and discourse-level patterns inherent in both visual data and language, as well as the many ways in which these aspects interact.
To ground our argument, we examine five seemingly distinct tasks, each of which we claim exemplifies this broader multimodal challenge. Our aim is not to provide a survey of relevant tasks, but rather to illustrate the general problem with a few representative tasks. The tasks we consider are Change Captioning Jhamtani and Berg-Kirkpatrick (2018), Video Question Answering (Video QA) Zeng et al. (2017); Bansal, Zhang, and Chellappa (2020), Video Captioning Yao et al. (2015), Movie Auto Audio Description (Movie Auto AD) Han et al. (2023b, a), and Visual Storytelling Huang et al. (2016). Figure 1 presents an illustration and a prototypical example for each of the five tasks, drawn from publicly available datasets. By analyzing them through a shared lens, we highlight common modeling and evaluation challenges that are often overlooked when tasks are considered in isolation. Yet, we recognize that these tasks also differ in crucial ways that stem from their distinct communicative intents and objectives. For instance, crafting a coherent and imaginative narrative in Visual Storytelling poses fundamentally different demands compared to pinpointing specific visual changes in Change Captioning. These task-specific nuances, which can be ignored or flattened when focusing solely on shared technical challenges, underscore the need for specific modeling proposals and evaluation protocols. This constitutes the second key argument of our position paper.
Our goal is to promote a rethinking of how language generation from visual events is conceptualized, studied, and evaluated within multimodal NLP. In doing so, we aim to (i) articulate a unified framework for understanding multimodal tasks involving visual event sequences, (ii) identify key open questions and methodological gaps, and (iii) propose directions for future research that draw on insights from human cognition and linguistic theory, recognizing the communicative specificities of each task. We contend that improving models’ ability to understand and generate language about visual events is both timely and necessary, both for practical applications and to advance our knowledge of how visual events are understood, interpreted, and narrated by humans and machines. The main contributions of this work are as follows:
  • We offer a comprehensive synthesis of modeling approaches, training paradigms, and evaluation methodologies used across five vision-to-language tasks involving visual events. By identifying shared pillars and methodological commonalities, we provide a unified lens that we hope will serve as a foundation for future research.
  • We identify and articulate a set of core challenges that recur across tasks dealing with temporally ordered visual events. These include tracking and grounding of entities over time, modeling event continuity and causality, and maintaining coherence and cohesion in the generated language.
  • We show that, within the problem of generating language from visual events, different tasks vary along various features based on their communicative intents. To this end, we provide a quantitative characterization of the tasks based on two dimensions—visual similarity and textual consistency—capturing the similarity of the elements in the visual input, and the consistency of the textual output.
  • We outline several promising research directions aimed at advancing models’ ability to capture the complex, time-dependent relationships between visual input and linguistic output. These include cognitively inspired representations, improved discourse modeling, and principled evaluation metrics that reflect narrative quality.
{tblr}
@@
Refer to caption The blue truck is no longer there. A car is approaching the parking lot from the right.
@@
Refer to caption The person is sitting at the dining room table wrapped in a blanket. The person is eating cereal and drinking orange juice.
@@
Refer to caption Sully adjusts his seat harness. A male passenger looks up from his magazine … Sully sticks out an arm as the jet bellies down onto the river.
<Write a story for this image sequence.,
@@
Refer to caption A discus got stuck up on the roof. Why not try getting it down with a soccer ball? … It didn’t work so we tried … are all stuck on the roof.
@@
2 Model Architectures
Modeling approaches to text generation from visual events have evolved over time from being based on recurrent neural networks (RNN; Hochreiter and Schmidhuber, 1997) to being transformer-based Vaswani et al. (2017). More recent models directly leverage pre-trained large (vision)-language models (LLMs/VLMs), often in a zero-shot manner. In this section, we discuss this evolution and summarize the various state-of-the-art model architectures proposed for this problem. We ground our discussion in five representative tasks—listed in Figure 1—that serve to exemplify the main trends. Change Captioning requires generating a caption to describe semantic changes between a pair of related input images. This task is particularly relevant for applications involving surveillance or quality inspection of ‘before’ and ‘after’ visual data. The goal of Visual Storytelling is to generate a fictitious, coherent narrative from a temporally-ordered sequence of images or frames. Applications of this task include digital storytelling and media creation. Video Captioning concerns describing videos by summarizing them, while Video QA involves generating appropriate answers to questions about the entities and events in videos. Both tasks have a wide range of applications, such as content summarization on streaming platforms and interacting with virtual assistants. Movie Auto AD involves generating narrative descriptions to complement the dialogues and soundtracks of movies. Enhancing the accessibility of movies and automatizing dubbing are some of the applications of this task.
@@ 2 Model Architectures
These tasks originate from varied research communities that focus on different modalities (language vs. vision) and place emphasis on different applications. However, as we argue in this section, it is possible to identify an underlying modeling approach common to tasks dealing with text generation from visual events. Architectures proposed for these tasks primarily comprise three modules—a vision encoder, a language decoder, and an intermediate module (typically referred to as the projector or adapter) for adapting visual information into contextualized representations for text generation. We describe these modules’ functionality and review the common design principles across tasks. Furthermore, we also discuss how off-the-shelf pre-trained vision-language models (VLMs) are currently being used to handle tasks involving text generation from visual events. We provide an overview of the models we reviewed for each task in Table 1.
@@ 2.1 Vision Encoder
The primary purpose of a vision encoder in vision-to-language tasks is to extract information from the visual input and to optimally encode it into a contextual representation that guides language generation. To achieve this, encoders in the proposed models follow multiple steps, some of which are common across language generation tasks from visual events. First, a pre-trained vision model is utilized for extracting feature representations of the raw input sequences of images or video frames. Earlier approaches used vision models based on convolutional neural networks (CNNs) such as ResNet He et al. (2016) or R3D Tran et al. (2018) that are primarily pre-trained on the object detection task using large amounts of image or video data. Most of the recent models across tasks use transformer-based vision models pre-trained for various image-only and image-text alignment objectives, e.g., CLIP-ViT-L Radford et al. (2021). We note that besides the primary input sequence of images or video frames, models proposed for some of the tasks, e.g., Movie Auto AD, utilize additional input data such as close-ups of characters in the movie clips (exemplars) Han et al. (2023a). Furthermore, the TAPM Yu et al. (2021) model proposed for the Visual Storytelling task utilizes FasterRCNN Ren et al. (2015) for extracting such local entity-level features alongside the global image-level features from ResNet.
Following the extraction of visual features using pre-trained vision models, most vision encoders comprise an internal sequence-encoder component for learning relationships and dependencies between the individual image or frame-level features at different temporal positions. Some models implement this step either using RNNs or a transformer network with multi-head self-attention for learning temporal relationships and position-encoding mechanism for tracking the order of entities or events in the visual sequence.
Beyond these common components across tasks, vision-encoders may also contain additional task-specific steps for capturing the visual information in a way that suits the task’s objective better. For instance, the ViLA model Wang et al. (2025) for the Video QA task utilizes a learnable Frame-Sampler component to efficiently select a small subset of frames that are most likely to contain the relevant information needed to answer the question. Another example with a task-specific step is the MSCM+BART model Hsu et al. (2020) for Visual Storytelling, in which the initial set of image objects or ‘concepts’ is expanded using an external knowledge graph for generating diverse and informative stories. Despite these task-specific steps, we found that the vision encoder module in the architectures proposed for the various language generation tasks dealing with visual events share a common set of components that are broadly outlined in Figure 2.
Figure 2: Outline of the architecture common across modeling approaches for tasks involving language generation from visual events.